OpenStudy (anonymous):

Solve x^2 + 18x - 40 = 0

OpenStudy (mertsj):

Can you factor it?

OpenStudy (anonymous):

What multiplies to make 40??

OpenStudy (anonymous):

what do you the answer is

OpenStudy (anonymous):

2 and 20 make 40, and as to the rest?

OpenStudy (anonymous):

(x plus or minus __)(x plus or minus __)

OpenStudy (mertsj):

(x+20)(x-2)=0

OpenStudy (mertsj):

x+20 = 0 or x-2 = 0

OpenStudy (anonymous):

set that equal to 0 separately

OpenStudy (mertsj):

x=-20 or x = 2